# Hyperuniformity and optimal transport of point processes

### v3

@@ 2 Main results
In this section, we state our main results. Firstly, we state our global matching result (Theorem 1) for stationary processes with integrable RPCM which shows that generic point processes in dimension $d\geq 3$ cannot have worser matching cost bounds than that of the Poisson process. Further, assuming hyperuniformity, we show in Theorem 2 that the bounds can be improved in dimensions $d=1,2$. We also discuss our results in context of existing results, methods and provide examples of point processes verifying the above bounds. The proofs of these results proceed via bounds for matching or transport cost between a point process and Lebesgue measure on large boxes. This is the content of Theorems 3 and 4, which are of independent interest and can also be viewed as an equivalent formulation of good transport between point processes. The proofs of all these main theorems - Theorems 1, 2, 3 and 4 - are deferred to Section 3. In Section 2.3, we comment on extension to random measures.
@@
Let $\mu,\nu$ independent simple stationary processes with unit intensity having integrable RPCM. There exists a translation-invariant matching $T$ between $\mu$ and $\nu$ such that the typical matching distance $X$ (see (2)) satisfies the following bound
@@
[…]
@@
The bounds for $d=1,2$ are following the methods of [30, Theorem 1(ii)] and are stated here for completeness. More precisely, doing arbitrary matching on an increasing sequence of dyadic cubes, one can show that there exists a matching from $\nu$ to $\mu$ such that
@@
for some finite $c$ and with $\sigma_{\cdot}(r)$ as defined in (4); see the proof of Theorem 1 for more details on this derivation.
Thus, in the integrable RPCM case, $\sigma$ is bounded and so we have that $\mathbf{E}(X^{\alpha})<\infty$ for all $\alpha<d/2$ and in $d=1,2$, we have $\mathbf{E}(w(X))<\infty$ for $w$ as above. To check this, verify by contradiction that for $x,r>10,x^{d/2}\ln(x)^{-\gamma}>r$ implies $x^{d/2}>cr\ln(r)^{\gamma}$. So, our Theorem 1 improves this bound in $d=3,4$ and in $d\geq 5$, [30, Theorem 1(ii)] (i.e., (7)) yields better bounds
We now turn to hyperuniform point processes, i.e. we assume $\sigma(r)\to 0$ as $r\to\infty$. From [46, Lemma 1.6] (see also [7, Theorem 2A] and [9]), we know that $\sigma(r)\geqslant Cr^{-1}$ for some $C>0$. Under the assumption of $|x||\beta(\mathrm{d}x)|$ being integrable, it is known that $\sigma(r)=\Theta(r^{-1})$ [45, Proposition 2]. Thus (7) can at the best yield
@@
    $\int_{\mathbb{R}^{2}}\mathbf{1}_{\{|x|>1\}}\ln(|x|)|\beta|(\mathrm{d}x)<\infty.$
@@
    The same results hold when $\mathbb{Z}^{2}$ are replaced by an independent stationary HU process $\nu$ satisfying the same conditions as $\mu$.
@@
    In $d=1$, suppose $\beta$ further satisfies $\int_{\mathbb{R}}|x|\,|\beta|(\mathrm{d}x)<\infty$ and let $\nu$ be an independent simple point process satisfying the same assumptions as $\mu$. Then there exists a translation invariant matching $T$ from $\mu$ to $\nu$ with typical distance $X$ such that
@@
Regarding the last statement in Item (1), it is simply obtained by buiding matchings $T:\mu\to\mathbb{Z}^{d},T^{\prime}:\nu\to\mathbb{Z}^{d}$ on a product probability space, then $T^{\prime}\circ T^{-1}$ is a matching between $\mu$ and $\nu$ with the same properties; in this regard the result involving $\mathbb{Z}^{d}$ is stronger, even though lattice or shifted lattice doesn’t have an integrable RPCM. It seems here that the independence assumption between $\mu$ and $\nu$ is superfluous but it is actually necessary because in the dependent case, the respective matchings of $\mu,\nu$ are built on a new probability space, and it is not clear they can be recombined on the same probability space by maintaining the same depedency. We also remark that the logarithmic integrability condition in Item (1) implies that $\ln(n)\,\sigma_{\mu}(n)=o(1)$ and the integrability condition in Item (2) implies that $n\sigma_{\mu}(n)=O(1)$ which in $d=1$, is same as boundedness of $\text{Var}(\mu(\Lambda_{n}))$.
@@
For the proofs of Theorems 1 (for $d\geq 3$) and 2, we derive transport rates (i.e., cost of an allocation scheme) between the point process restricted to a finite box and the restriction of the Lebesgue measure to that box using Fourier-analytic bounds due to Bobkov and Ledoux [13]; see Theorems 3 and 4. More specifically, consider the restricted sample $\mu_{n}=\mu 1_{\Lambda_{n}}$, where $\Lambda_{n}=[-\pi n^{1/d},\pi n^{1/d}]^{d}$ , and its renormalisation $\tilde{\mu}_{n}=\frac{n}{N}\mu_{n}$ (where $N=\mu(\Lambda_{n})$), for some ‘nice’ cost function $w$, to obtain the linear cost
@@
where $\widetilde{\mathsf{C}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n})$ is the cost with respect to $w$ of the optimal matching (or transport or allocation scheme) under the Toroidal metric on $\Lambda_{n}$ between $\tilde{\mu}_{n}$ and ~$\tilde{\mathcal{L}}_{n}$, the Lebesgue measure on $\Lambda_{n}$ normalized to have total measure $n$; see Section 2.1 for precise definitions. A similar bound is obtained for transport rates between $\tilde{\nu}_{n}$ (a suitably normalized version of $\nu$ in $\Lambda_{n}$) and ~$\tilde{\mathcal{L}}_{n}$. Using triangle inequality and Birkhoff-von Neumann theorem, one can obtain a matching between $\tilde{\mu}_{n}$ and $\tilde{\nu}_{n}$ which is then extended to the infinite space via relative compactness arguments and that $\tilde{\mu}_{n},\tilde{\nu}_{n}$ converge to $\mu,\nu$ in vague topology. This is done in Proposition 11 in Section 3.3, thereby yielding that there is a matching $T:\nu\to\mu$ of the limiting infinite measures such that
@@
which indeed corresponds to the definition of the typical distance (2).
In $d=1,2$, we bound $n^{-1}\mathbf{E}\widetilde{\mathsf{C}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n})$ by powers of $\ln(n)$ and though these indicate bounds as in Theorem 1 for $d=1,2$ but we are unable to deduce them from this. As indicated in Remark 1, we use the methods from [30] for $d=1,2$. However, we mention that our finite volume bounds in Theorem 3 match with those for i.i.d. points in all dimensions and also applies to non-integrable RPCM case. The same comment also applies to the case $d=1$ in Theorem 2.
@@
We now comment briefly on the differences of our approach with respect to other results on hyperuniform processes. Both in the specific examples such as [53, 35, 49] and the general results of [17] the authors use also strong moment assumptions on the point process. The proof of [53] proceeds by constructing a Whitney-type partition with respect to a modified metric and a potential-theoretic lemma; see also [54]. [35] also uses potential-theoretic estimates. Such potential-theoretic estimates work well for zeros of Gaussian entire functions or Ginibre ensemble. [49, 17] rely upon sub-additivity with suitable Whitney-type decompostions and the crucial estimates of the error in sub-additive arguments are achieved using PDE or functional-analytic tools as for combinatorial optimization problems over i.i.d. points in [3, 25, 26]. Both the potential-theoretic or PDE approach involve estimating solutions of a certain distributional Poisson equation. Our approach stemming from [12] is very different to these and involves only estimates based on the pair correlation function of the point processes. While the proof strategy is relatively simpler and works in greater generality, our methods cannot give higher-moment estimates unlike in the above articles (when the assumptions are sufficiently strong). It is important to mention that [32] also recovers Theorem 2 by showing that a hyperuniform process for which the normalized variance (i.e., (4)) goes to 00 at a logarithmic speed, the Coulomb energy is finite, and the latter implies finite mean $2$-Wasserstein distance to Lebesgue measure (or equivalently to the lattice); see Remark 1.8 therein. Interestingly, an analogous result to Theorem 2 in the deterministic setting which is existence of bounded Lipschitz matching, was recently proven in [40] under a similar logarithmic condition on discrepancy and also using Fourier-analytic methods.
@@ 2.1 Optimal transport of large samples
By [50, Theorem 1.7], if $\mu,\nu$ have same finite total mass and $c$ is lower semi-continuous then the infimum is indeed reached by some $M$, called the optimal transport plan or coupling. Since we will work with continuous cost functions (modulus of continuity or $x^{2}$) existence of optimal coupling is always guaranteed for us. If furthermore $\mu,\nu$ are supported on some cube $\Lambda$, let $d_{\Lambda}(x,y)$ be the toric distance on $\Lambda$, and the corresponding cost is
$\widetilde{\mathsf{C}}_{w}(\mu,\nu)=\widetilde{\mathsf{C}}_{w}(\mu,\nu;\Lambda):=\inf_{M}\int_{\Lambda^{2}}w(d_{\Lambda}(x,y))M(\mathrm{d}x,\mathrm{d}y).$
If $\mu=\nu=0$ (i.e., the null measure), by definition $\mathsf{C}_{w}(\mu,\nu)=\widetilde{\mathsf{C}}_{w}(\mu,\nu)=0$. We take the toric distance as our main tool relies on Fourier decomposition on the torus, but in the large window asymptotics, transport properties for toric and Euclidean distance will coincide. We often drop the dependence on domain $\Lambda$ when it is obvious.
@@
Our standard assumption in Theorems 1 and 2 is about an integrable RPCM $\beta$. However in our upcoming finite sample bounds, we shall explicitly quantify the dependence of matching or transport cost on the tail of $\beta$, which might not be integrable. We now introduce two quantities measuring the tail of $\beta$. Define
$\mathsf{b}_{n}:=1+\int_{\Lambda_{2^{d}n}}|\beta|(\mathrm{d}x),$ (8)
@@
Note that under integrability of $\beta$, $\mathsf{b}_{n}$ is a bounded sequence and $\varepsilon(t)$ appears implicitly in the logarithmic integrability condition in Theorem 2.
@@
More formally, we consider independent stationary point processes $\mu,\nu$, and upper bound their transport cost on the large cube $\Lambda_{n}=[-\pi n^{1/d},\pi n^{1/d}]^{d}.$ Using the Fourier-analytic method, which goes through estimating Fourier-Stieljes transform of the restricted samples $\tilde{\mu}_{n}:=\frac{n}{N}\mu 1_{\Lambda_{n}}$, where $N=\mu(\Lambda_{n})$ is the number of points, we obtain the following bounds.
@@
Let $\mu$ be a stationary point process with unit intensity. Denote by $\tilde{\mu}_{n}=\frac{n}{N}\mu 1_{\Lambda_{n}}$ and ~$\tilde{\mathcal{L}}_{n}:=\frac{1}{(2\pi)^{d}}\mathcal{L}_{n}$, the renormalised samples with mass $n$ on $\Lambda_{n}.$ When $N=0$, set $\tilde{\mu}_{n}=n\delta_{0}$. Then we have the following bounds on the squared Wasserstein costs
@@
[…]
@@
    $\alpha_{2}(n)=c\,n\,\big{(}1+\int_{n^{-1/d}}^{c_{0}}\varepsilon(r)\,r^{d-3}\,{\mathrm{d}r}\big{)}.$ (15)
@@
    $\int_{\mathbb{R}^{2}}\mathbf{1}_{\{|x|>1\}}\,\ln(|x|)\,|\beta|(\mathrm{d}x)<\infty,$
@@
  3. 3.
    In $d=1$ if $\int_{\mathbb{R}}|x|\,|\beta|(\mathrm{d}x)<\infty$ then we have that $\mathbf{E}{\widetilde{\mathsf{W}}}_{2}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\,\leqslant c\,n\,\ln(n).$
The proofs are in Section 3.2. The choice $\tilde{\mu}_{n}=n\delta_{0}$ for $N=0$ is in some sense the worst choice but for our results one could have chosen any other measure with mass $n$. Note the linear rate of growth for $W_{2}^{2}$-transport cost in Theorem 3 for $d\geq 3$ and in Theorem 4(2) for $d=2$ respectively. These will be crucial to the proofs of Theorem 1 and 2. Though we cannot obtain linear-rate of growth in $d=1$ for HU processes, we have improved the bounds from Theorem 3.
Theorem 3 for Wasserstein distance is consistent with those known for i.i.d. points or mixing point processes or point processes generated via mixing Markov chains; see for example [23, Section 7], [13, Section 5], [14, Section 2], [15, Theorem 2] and [19, Proposition B.1]. Like with all our other main theorems, there are two significant differences. Firstly, our bound is for random points and for general point processes, no mixing-type assumption is required. In case the point processes have integrable reduced pair correlations (which is only about the second order marginal and weaker than the mixing assumptions in the cited articles), we recover the same bound as [13, 15, 23] for $2$-Wasserstein distance. Secondly in Theorem 4, we exhibit the particular behaviour of HU processes whereas the above papers do not distinguish between hyperuniform and non-hyperuniform point processes as this concept is not related to mixing.
As noted earlier, the above two theorems rely on Fourier-analytic bounds of Bobkov and Ledoux [12, 13, 14] and as in their works, it is possible to state the finite-volume bounds for more general costs even though the analogue in infinite-volume is unclear. Consider general $w$-transport costs $\tilde{C}_{w}$ where $w:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a continuity modulus, i.e. non-decreasing, continuous and sub-additive ($w(x+y)\leq w(x)+w(y)$). Observe that $w(x)=x^{p}$, $p\in[0,1]$ are continuity moduli. Though we use only these choices in our results, for completeness we state the general bound here.
@@
$\mathbf{E}\widetilde{{\mathsf{C}}}_{w}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leq c\,n\,\mathsf{b}_{n}\,\sum_{k=1}^{t_{0}}k^{d-1}q(k)\,w(n^{1/d}k^{-1})^{2}+cn^{2}w(n^{1/d}t_{0}^{-1})^{2},$
@@
(see (38)) and in particular, with the triangle inequality, $\text{Var}(N)\leqslant cn\varepsilon(n^{-1/d}))$ for some $c<\infty$. If the reverse inequality (i.e., $\text{Var}(N)\geqslant cn\varepsilon(n^{-1/d}))$) were to hold as one would expect in many cases, it would allow us to equivalently state our main theorems in terms of $\text{Var}(N)$ instead of $\beta$. However, the possibility of oscillating $\beta$ prevents us from doing so in the general case. Under the assumption that $\beta$ has negative sign, we have indeed
$\textrm{Var}(N)\asymp n\varepsilon(n^{-1/d}).$
This assumption means that in some sense the $2$-point correlation function is repulsive. Some determinantal point processes, or more generally, the so-called repulsive point processes satisfy this assumption because their $n$-point correlation function is repulsive at all orders (and everywhere); see [11].

### v4

@@ 2 Main results
In this section, we state our main results. Firstly, we state our global matching result (Theorem 1) for stationary processes with integrable RPCM which shows that generic point processes in dimension $d\geq 3$ cannot have worser matching cost bounds than that of the Poisson process. Further, assuming hyperuniformity, we show in Theorem 2 that the bounds can be improved in dimensions $d=1,2$. We also discuss our results in the context of other results, outline our proof ideas and provide examples of point processes verifying the above bounds. The proofs proceed via bounds for matching or transport cost between a point process and Lebesgue measure on large boxes. This is the content of Theorems 3 and 4, which are of independent interest and can also be viewed as an equivalent formulation of good transport between point processes. The proofs of all these main theorems - Theorems 1, 2, 3 and 4 - are deferred to Section 3. In Section 2.3, we comment on extension to random measures.
@@
Let $\mu,\nu$ be independent simple stationary processes with unit intensity and having integrable RPCM. There exists a translation-invariant matching $T$ between $\mu$ and $\nu$ such that the typical matching distance $X$ (see (4)) satisfies the following bound
@@
$w(x)=\begin{cases}\frac{\sqrt{x}}{1+|\ln(x)|^{\gamma}}$ if $d=1\\ \frac{x}{1+|\ln(x)|^{\gamma}}$ if $d=2\\ x^{2}$ if $d\geqslant 3.\end{cases}$
@@
The bounds for $d=1,2$ are obtained following the methods of [33, Theorem 1(ii)] and are stated here for completeness. We can show that there exists a matching from $\nu$ to $\mu$ whose typical distance $X$ satisfies
@@
for some finite $c$ and with $\sigma_{\cdot}(r)$ as defined in (6); see the proof of Theorem 1 at Section 3.4 for more details on this derivation.
Thus, in the integrable RPCM case, $\sigma$ is bounded and so we have that $\mathbf{E}(X^{\alpha})<\infty$ for all $\alpha<d/2$ and in $d=1,2$, we have $\mathbf{E}(w(X))<\infty$ for $w$ as above. So, our Theorem 1 improves this bound in $d=3,4$ and in $d\geq 5$, [33, Theorem 1(ii)] (i.e., (9)) yields better bounds.
We now turn to hyperuniform point processes, i.e. we assume $\sigma(r)\to 0$ as $r\to\infty$. From [50, Lemma 1.6] (see also [8, Theorem 2A] and [10]), we know that $\sigma(r)\geqslant Cr^{-1}$ for some $C>0$. Under the assumption of $|x||\beta(\mathrm{d}x)|$ being integrable, it is known that $\sigma(r)=\Theta(r^{-1})$ [48, Proposition 2]. Thus (9) can at the best yield
@@
    $\int_{\mathbb{R}^{2}}\mathbf{1}[|x|>1]\ln(|x|)|\beta|(\mathrm{d}x)<\infty.$
@@
    The same results hold when $\mathbb{Z}^{2}$ is replaced by an independent stationary HU process $\nu$ satisfying the same conditions as $\mu$.
@@
    In $d=1$, suppose the number variance $\textrm{Var}\left(\mu(\Lambda_{n})\right),n\geq 1$ is uniformly bounded and let $\nu$ be an independent simple point process satisfying the same assumptions as $\mu$. Then there exists a translation invariant matching $T$ from $\mu$ to $\nu$ with typical distance $X$ such that
@@
Remark 2.
Like for Theorem 1, the proof for $d=1$ borrows the approach of [33]; see Remark 1.
Regarding the last statement in Item (1), it can be obtained by building matchings $T:\mathbb{Z}^{2}\to\mu,T^{\prime}:\mathbb{Z}^{2}\to\nu$ on a product probability space, then $T\circ(T^{\prime})^{-1}$ is a matching between $\mu$ and $\nu$ with the same properties; in this regard the result involving $\mathbb{Z}^{2}$ is stronger, even though lattice or shifted lattice doesn’t have an integrable RPCM. However, we first prove the bounds in the case of independent point processes with integrable RPCM and then use our toy example of a cloaked perturbed lattice from Example 1 to obtain an invariant matching from $\mathbb{Z}^{2}$. It seems here that the independence assumption between $\mu$ and $\nu$ is superfluous but in the dependent case, the respective matchings of $\mu,\nu$ are built on a new probability space, and it is not clear how they can be recombined on the same probability space by maintaining the same dependency. We also remark that the logarithmic integrability condition in Item (1) implies that $\ln(n)\,\sigma_{\mu}(n)=o(1)$.
Regarding point (2), boundedness of the variance in dimension $1$ holds true in several examples; see for instance [1] who study models where this assumption is satisfied and give several examples, such as the 1D log gas.
@@
For the proofs of Theorems 1 (for $d\geq 3$) and 2, we derive transport rates (i.e., cost of an allocation scheme) between the point process restricted to a finite box and the restriction of the Lebesgue measure to that box using Fourier-analytic bounds due to Bobkov and Ledoux [14]; see Theorems 3 and 4. More specifically, consider the restricted sample $\mu_{n}=\mu\mathbf{1}_{\Lambda_{n}}$, where $\Lambda_{n}=(-\pi n^{1/d},\pi n^{1/d}]^{d}$ , and its renormalisation $\tilde{\mu}_{n}=\frac{n}{N}\mu_{n}$ (where $N=\mu(\Lambda_{n})$), for some ‘nice’ cost function $w$. One obtains the linear cost
@@
where $\widetilde{\mathsf{C}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n})$ is the cost with respect to $w$ of the optimal transport or allocation scheme under the Toroidal metric on $\Lambda_{n}$ between $\tilde{\mu}_{n}$ and $\tilde{\mathcal{L}}_{n}=\mathcal{L}(\Lambda_{n})^{-1}\mathcal{L}\mathbf{1}_{\Lambda_{n}}$; see Section 2.1 for precise definitions. A similar bound is obtained for transport rates between $\tilde{\nu}_{n}$ (a suitably normalized version of $\nu$ in $\Lambda_{n}$ with total mass $n$) and $\tilde{\mathcal{L}}_{n}$. Using triangle inequality, one can obtain a coupling between $\tilde{\mu}_{n}$ and $\tilde{\nu}_{n}$ which is then extended to the infinite space via relative compactness arguments and such that $\tilde{\mu}_{n},\tilde{\nu}_{n}$ converge to $\mu,\nu$ in the vague topology. Finally, using a result from [25], we show that the coupling between $\mu,\nu$ is realized by a matching. This is done in Proposition 11 in Section 3.3, thereby yielding that there is a matching $T:\nu\to\mu$ of the limiting infinite measures such that
@@
which indeed corresponds to the definition of the typical distance (3). We allow for the possibility that the matching and the point processes $\mu,\nu$ may need to be defined on a new probability space.
In dimensions $d=1,2$, the bounds on $n^{-1}\mathbf{E}\widetilde{\mathsf{C}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n})$ from Theorem 3 below are diverging and do not allow us to derive the result of Theorem 1, which is why we use the method of [33]. However, we mention that our finite volume bounds in Theorem 3 match with those for i.i.d. points in all dimensions and also applies to the non-integrable RPCM case. The same comment would also apply to the case $d=1$ in the setting of Theorem 4.
@@
We now comment on our approach in the context of other matching or transport results for hyperuniform processes. Both in the specific examples such as [56, 38, 52] and the general results of [18] the authors use also strong moment assumptions on the point process. The proof of [56] proceeds by constructing a Whitney-type partition with respect to a modified metric and a potential-theoretic lemma; see also [57]. [38] also uses potential-theoretic estimates. Such potential-theoretic estimates work well for zeros of Gaussian entire functions or Ginibre ensemble. [52, 18] rely upon sub-additivity with suitable Whitney-type decompostions and the crucial estimates of the error in sub-additive arguments are achieved using PDE or functional-analytic tools as for combinatorial optimization problems over i.i.d. points in [4, 28, 29]. Both the potential-theoretic or PDE approach involve estimating solutions of a certain distributional Poisson equation. Our approach stemming from [13] is very different to these and involves only estimates based on the pair correlation function of the point processes. While the proof strategy is relatively simpler and works in greater generality, our methods cannot give higher-moment estimates unlike in the above articles (when the assumptions are sufficiently strong). It is important to mention that [35] also recovers Theorem 2 by showing that a hyperuniform process for which the normalized variance (i.e., (6)) goes to 0 at a logarithmic speed, the Coulomb energy is finite, and the latter implies finite mean $2$-Wasserstein distance to Lebesgue measure (or equivalently to the lattice); see Remark 1.8 therein.
Interestingly, an analogous result to Theorem 2 in the deterministic setting which is existence of bounded Lipschitz matching, was recently proven in [42] under a similar logarithmic condition on discrepancy and also using Fourier-analytic methods. Also in specific examples, it is quite likely that the matching bounds can be improved significantly; for example, very recently [24] have proven that perturbed lattices satisfy matching tail bounds corresponding to their hole probabilities.
@@ 2.1 Optimal transport of large samples
By [53, Theorem 1.7], if $\mu,\nu$ have same finite total mass and $c$ is lower semi-continuous then the infimum is indeed reached by some $M$, called the optimal transport plan or coupling. Since we will work with continuous cost functions (modulus of continuity or $x^{2}$) existence of optimal coupling is always guaranteed for us. If furthermore $\mu,\nu$ are supported on some cube $\Lambda$, let $d_{\Lambda}(x,y)$ be the toric distance on $\Lambda$, and the corresponding cost is
${\widetilde{\mathsf{C}}_{w}(\mu,\nu)}=\widetilde{\mathsf{C}}_{w}(\mu,\nu;{\Lambda}):=\inf_{M}\int_{\Lambda^{2}}w(d_{\Lambda}(x,y))\,M(\mathrm{d}x,\mathrm{d}y).$ (10)
If $\mu=\nu=0$ (i.e., the null measure), by definition $\mathsf{C}_{w}(\mu,\nu)=\widetilde{\mathsf{C}}_{w}(\mu,\nu)=0$. We take the toric distance as we rely on Fourier decomposition on the torus, but in the large window asymptotics, the mean transport cost for the toric and Euclidean metric coincide under our assumptions. We often drop the dependency on the domain $\Lambda$ when it is obvious.
@@
Our standard assumption in Theorems 1 and 2 is about an integrable RPCM $\beta$. However in our upcoming finite sample bounds, we shall explicitly quantify the dependence of matching or transport cost on the tail of $\beta$, which might not be integrable. We now introduce two quantities measuring the tail of $\beta$. Recalling $\Lambda_{n}=(-\pi n^{1/d},\pi n^{1/d}]^{d},$ define
$\mathsf{b}_{n}:=1+|\beta|(\Lambda_{n}+\Lambda_{n})=1+\int_{\Lambda_{2^{d}n}}|\beta|(\mathrm{d}x),$ (11)
@@
Note that under integrability of $\beta$, $\mathsf{b}_{n}$ is a bounded sequence and $\varepsilon(t)$ appears implicitly in the logarithmic integrability condition in Theorem 2 and upcoming Theorem 4(2).
@@
More formally, we consider independent stationary point processes $\mu,\nu$, and upper bound their transport cost on the large cube $\Lambda_{n}$. Using the Fourier-analytic method, which goes through estimating Fourier-Stieljes transform of the restricted samples $\tilde{\mu}_{n}:=\frac{n}{N}\mu\mathbf{1}_{\Lambda_{n}}$, where $N=\mu(\Lambda_{n})$ is the number of points, we obtain the following bounds.
@@
Let $\mu$ be a stationary point process with unit intensity. Denote by $\tilde{\mu}_{n}=\frac{n}{N}\mu\mathbf{1}_{\Lambda_{n}}$ and $\tilde{\mathcal{L}}_{n}:=\frac{1}{(2\pi)^{d}}\mathcal{L}\mathbf{1}_{\Lambda_{n}}$, the renormalised samples with mass $n$ on $\Lambda_{n}.$ When $N=0$, set $\tilde{\mu}_{n}=n\delta_{0}$. Then we have the following bounds on the squared Wasserstein costs
@@
$\alpha_{2}(n)\,:=\,$ ${\mathsf{b}_{n}}\times\begin{cases}n^{2}$ if $d=1\\ n\ln(n)$ if $d=2\\ n$ if $d\geqslant 3,\end{cases}$ (15)
@@
    $\alpha_{2}(n):=c\,n\,\big(1+\int_{n^{-1/d}}^{c_{0}}\varepsilon(r)\,r^{d-3}\,{\mathrm{d}r}\big).$ (18)
@@
    $\int_{\mathbb{R}^{2}}\mathbf{1}[|x|>1]\,\ln(|x|)\,|\beta|(\mathrm{d}x)<\infty,$
@@
The proofs are in Section 3.2. The choice $\tilde{\mu}_{n}=n\delta_{0}$ for $N=0$ is in some sense the worst choice but for our results one could have chosen any other measure with mass $n$. Note the linear rate of growth for $W_{2}^{2}$-transport cost in Theorem 3 for $d\geq 3$ and in Theorem 4(2) for $d=2$ respectively. These will be crucial to the proofs of Theorem 1 and 2. Though we cannot obtain linear-rate of growth in $d=1$ for HU processes, we could improve the bounds from Theorem 3 as follows. If $\int_{\mathbb{R}}|x|\,|\beta|(\mathrm{d}x)<\infty$ then we have that $\mathbf{E}{\widetilde{\mathsf{W}}}_{2}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};{\Lambda_{n}})\,\leqslant c\,n\,\ln(n).$ Since this is not used in our main results, we do not provide a proof.
Theorem 3 for Wasserstein distance is consistent with those known for i.i.d. points or mixing point processes or point processes generated via mixing Markov chains; see for example [26, Section 7], [14, Section 5], [15, Section 2], [16, Theorem 2] and [20, Proposition B.1]. Like with all our other main theorems, there are two significant differences. Firstly, our bound is for random points and for general point processes, no mixing-type assumption is required. In case the point processes have integrable reduced pair correlations (which is only about the second order marginal and weaker than the mixing assumptions in the cited articles), we recover the same bound as [14, 16, 26] for $2$-Wasserstein distance. Secondly in Theorem 4, we exhibit the particular behaviour of HU processes whereas the above papers do not distinguish between hyperuniform and non-hyperuniform point processes as this concept is not related to mixing.
As noted earlier, the above two theorems rely on the idea of Bobkov and Ledoux to combine Fourier-analytic bounds with a smooth convolution [13, 14, 15] and as in their works, it is possible to state the finite-volume bounds for more general costs even though the analogue in infinite-volume is unclear. Consider general $w$-transport costs $\tilde{C}_{w}$ where $w:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a continuity modulus, i.e. non-decreasing, continuous and sub-additive ($w(x+y)\leq w(x)+w(y)$). Observe that $w(x)=x^{p}$, $p\in[0,1]$ are continuity moduli. Though we use only these choices in our results, for completeness we state the general bound here.
@@
$\mathbf{E}\widetilde{{\mathsf{C}}}_{w}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leq c\,n\,\mathsf{b}_{n}\,\sum_{k=1}^{t_{0}}k^{d-1}q(k)\,w(n^{1/d}k^{-1})^{2}+cn^{2}w(n^{1/d}t_{0}^{-1})^{2}.$
@@
(see (43)) and in particular, with the triangle inequality, $\text{Var}(N)\leqslant cn\varepsilon(n^{-1/d}))$ for some $c<\infty$. If the reverse inequality (i.e., $\text{Var}(N)\geqslant cn\varepsilon(n^{-1/d}))$) were to hold as one would expect in many cases, it would allow us to equivalently state our main theorems in terms of $\text{Var}(N)$ instead of $\beta$. However, the possibility of oscillating $\beta$ prevents us from doing so in the general case. Under the assumption that $\beta$ has negative sign, we show in Lemma 8 that
$\textrm{Var}(N)\asymp n\varepsilon(n^{-1/d});$
The negative sign assumption means that in some sense the $2$-point correlation function is repulsive. Some determinantal point processes, or more generally, the so-called repulsive point processes satisfy this assumption because their $n$-point correlation function is repulsive at all orders (and everywhere); see [12].